# Deep Unfolding Network with Spatial Alignment for multi-modal MRI reconstruction

### v1

@@ 3.3 Network Parameters and Loss Function
where $N$ represents the number of training samples, ${x}_{\text{rec},i}$ denotes the $i^{th}$ image reconstructed by the network, and ${x}_{\text{gt},i}$ signifies the $i^{th}$ ground-truth image.
@@ 4.1 Datasets
We evaluate our method using three datasets, namely the fastMRI dataset, the IXI dataset and an in-house dataset.
FastMRI Dataset [^1] https://fastMRI.med.nyu.edu/.: To maintain consistency in the experiments, we follow the setups described in (Xuan et al., 2022) and select 340 pairs of T1-weighted and T2-weighted axial brain MRIs. We use 170 volumes (2720 pairs of slices) as the training set, 68 volumes (1088 pairs of slices) as the validation set, and 102 volumes (1632 pairs of slices) as the test set. The in-plane size of all T1-weighted images and T2-weighted images is $320\times 320$, with the resolution of $0.68mm\times 0.68mm$ and the slice spacing of 5mm.
IXI Dataset [^2] http://brain-development.org/ixi-dataset/.: The IXI dataset contains 576 paired multi-modal 3D brain MRIs. We select 570 pairs of PD-weighted and T2-weighted axial brain MRIs for our experiments. Specifically, we use 285 volumes as the training set, 115 volumes as the validation set, and 170 volumes as the test set. The in-plane size of all PD-weighted images and T2-weighted images is $256\times 256$. Consistent with (Lei et al., 2023), we select the middle 100 slices from each volume for our experiments.
@@ 4.2 Compared Methods
To evaluate the performance of our model (DUN-SA), we conduct comparisons with state of the art methods, including two single-modal MRI reconstruction methods: E2E-Varnet (Sriram et al., 2020), HQS-Unet (Xin et al., 2022); and four multi-modal MRI reconstruction methods: MD-DUN (Yang et al., 2020b), MM-E2E-Varnet(a variant of E2E-VarNet for multi-modal input), SAN (Xuan et al., 2022), and MC-CDic (Lei et al., 2023). The details are presented as follows.
E2E-Varnet employs variational techniques to solve the MRI reconstruction problem, unfolding it into a network structure combined with U-Net for end-to-end learning.
4.3 Performance Evaluation Metric
To accurately and comprehensively evaluate the performance of image reconstruction, we select three core metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Metric (SSIM), and Mean Absolute Error (MAE). These metrics are employed to quantify the similarity between the reconstructed images and the corresponding ground truth. It is noteworthy that higher values of PSNR and SSIM, along with lower MAE values, indicate superior reconstruction performance of the model.
4.4 Implementation Details
In our experiments, we follow the paradigm set by the fastMRI challenge, where under-sampled MRIs are obtained by masking the corresponding fully-sampled MRIs in k-space using a Cartesian sampling pattern. Specifically, we employ two types of sampling patterns: random and equispaced. These are applied at sampling ratios of 25% (4x acceleration) and 12.5% (8x acceleration), respectively. Acknowledging that low-frequency signals contain the majority of energy in k-space, we allocate 32% of the sampling to these low frequencies in both patterns. The remaining portion of the sampling is distributed either randomly or in an equispaced manner.
All network models used in these experiments are implemented using the PyTorch framework and are run on a computer equipped with four NVIDIA GeForce GTX 3090 GPUs. During the model training phase, we set the batch size to be $2$, adopt an end-to-end training strategy, and use the SSIM loss function with a coefficient of $1$. All parameters are optimized using the Adam optimizer with a learning rate of $1\times 10^{-4}$. To prevent overfitting, we used a validation set to select the optimal parameters during the training process. The final experimental results are based on these optimal parameters, as selected and validated through the validation set, and applied on the test set.
Table 1: Quantitative evaluation of DUN-SA vs. other methods on the fastMRI dataset for 4x and 8x acceleration under equispaced 1D subsampling masks, where T1-weighted images are used as reference modality to assist the reconstruction of T2-weighted images. Best results are emphasized in bold, and the second best are emphasized with an underline.
Methods Equispaced 4x acceleration Equispaced 8x acceleration
@@ 4.4 Implementation Details
Table 2: Quantitative evaluation of DUN-SA vs. other methods on the fastMRI dataset for 4x and 8x acceleration under random 1D subsampling masks, where T1-weighted images are used as reference modality to assist the reconstruction of T2-weighted images. Best results are emphasized in bold, and the second best are emphasized with an underline.
Methods Random 4x acceleration Random 8x acceleration
PSNR SSIM MAE PSNR SSIM MAE
Zero-filling 27.12$\pm$1.03 0.7373$\pm$0.0282 0.0233$\pm$0.0032 23.88$\pm$1.70 0.5818$\pm$0.0431 0.0342$\pm$0.0067
@@
Fig. 4: Visual comparison with representative methods for 4x acceleration under 1D equispaced subsampling mask on fast MRI dataset. First row: Reconstructed images by different methods; second row: Zoomed-in region of interest; third row: Equispaced mask of 4x acceleration and error maps of different methods.
@@ 5.1 Results on the fastMRI dataset
Table 3: Quantitative evaluation of DUN-SA vs. other methods on the IXI dataset for 4x and 8x acceleration under equispaced 1D subsampling masks, where PD-weighted images are used as reference modality to assist the reconstruction of T2-weighted images. Best results are emphasized in bold, and the second best are emphasized with an underline.
Methods Equispaced 4x acceleration Equispaced 8x acceleration
@@
Table 4: Quantitative evaluation of DUN-SA vs. other methods on the IXI dataset for 4x and 8x acceleration under random 1D subsampling masks, where PD-weighted images are used as reference modality to assist the reconstruction of T2-weighted images. Best results are emphasized in bold, and the second best are emphasized with an underline.
Methods Random 4x acceleration Random 8x acceleration
PSNR SSIM MAE PSNR SSIM MAE
Zero-filling 26.32$\pm$2.04 0.6177$\pm$0.0465 0.0296$\pm$0.0068 23.91$\pm$2.01 0.5371$\pm$0.0538 0.0380$\pm$0.0086
@@
Fig. 5: Visual comparison with representative methods for 8x acceleration under 1D equispaced subsampling mask on the IXI dataset. First row: Reconstructed images by different methods; second row: Zoomed-in region of interest; third row: Equispaced mask of 8x acceleration and error maps of different methods.
We first run our proposed method on the fastMRI dataset. In this part of the experiments, we use the fully-sampled T1-weighted image as the reference modality to assist the k-space undersampled data of the T2 modality in reconstruction.
Table 1 and Table 2 display the quantitative evaluations on the fastMRI dataset for 1D equispaced and random masks under 4× and 8× acceleration. We observe that images reconstructed by multi-modal methods show significant improvements in the PSNR, SSIM, and MAE metrics across all experimental setups compared to those reconstructed by single-modal methods. This enhancement is particularly pronounced for 8x acceleration. The introduction of spatial alignment in the multi-modal reconstruction further improves the PSNR, SSIM, and MAE values. It is evident that DUN-SA outperforms all others in every metric across all settings. In comparison with the second-best values, for 8x acceleration under random 1D subsampling masks, the highest differences in PSNR, SSIM, and MAE are 1.38db, 0.0046, and 0.001, respectively. In contrast, for 4x acceleration under equispaced 1D subsampling masks, the smallest differences are 0.76db, 0.0019, and 0.0002. These results suggest that DUN-SA can effectively utilize the reference modality to aid the MRI reconstruction of the target modality, leading to a noticeable improvement in reconstruction quality.
Fig. 4 illustrates a qualitative comparison of the reconstructed images and their corresponding error maps using various methods on the fastMRI dataset for 4x acceleration under 1D equispaced mask. The first row displays the reconstruction resluts from different methods, the second row showcases zoomed-in regions of interest, and the third row demonstrates the error maps. The patterns in the error maps represent reconstruction errors; a less intense color on the map (closer to the cool end of the spectrum) signifies superior reconstruction quality. Clearly, the Zero-filling image exhibits aliasing artifacts and lacks anatomical details. Significantly, compared to other methods, our method reconstructs images with minimal visible artifacts and reconstruction discrepancies compared to the ground truth. Specifically, as depicted in the zoomed-in regions of interest in the second row, the detail indicated by the green arrow is only recovered by DUN-SA and MC-CDic. Moreover, at the location indicated by the red arrow, MC-CDic reconstruct a texture that originally does not exist, introducing a mistake. These findings verify that our method not only shows remarkable enhancements in all metrics but also reconstructs more textural details.
5.2 Results on the IXI dataset
Table 5: Quantitative evaluation of DUN-SA vs. other methods on the in-house dataset for 4x and 8x acceleration under equispaced 1D subsampling masks, where T1-weighted images are used as reference modality to assist the reconstruction of T2-weighted images. Best results are emphasized in bold, and the second best are emphasized with an underline.
Methods Equispaced 4x acceleration Equispaced 8x acceleration
@@ 5.2 Results on the IXI dataset
Table 6: Quantitative evaluation of DUN-SA vs. other methods on the in-house dataset for 4x and 8x acceleration under random 1D subsampling masks, where T1-weighted images are used as reference modality to assist the reconstruction of T2-weighted images. Best results are emphasized in bold, and the second best are emphasized with an underline.
Methods Random 4x acceleration Random 8x acceleration
PSNR SSIM MAE PSNR SSIM MAE
Zero-filling 26.82$\pm$0.66 0.7075$\pm$0.0144 0.0261$\pm$0.0024 24.27$\pm$0.66 0.5982$\pm$0.0164 0.0359$\pm$0.0033
@@
Fig. 6: Visual comparison with representative methods for 8x acceleration under 1D equispaced subsampling mask on the in-house dataset. First row: Reconstructed images by different methods; second row: Zoomed-in region of interest; third row: Equispaced mask of 8x acceleration and error maps of different methods.
@@
We validate the effectiveness of our method on a larger dataset, IXI. In this part of the experiments, the fully sampled PD-weighted image is used as reference modality image to assist the reconstruction of the paired undersampled T2-weighted image.
@@
Table 3 and Table 4 present the quantitative evaluations on the IXI dataset for 1D equispaced and random masks under 4× and 8× acceleration. From the metrics, it can be observed that the reference modality can significantly enhance the reconstruction performance, and the spatial alignment further improves all metrics, which is consistent with the experimental results on the fastMRI dataset. Our method once again brings noticeable improvements for every metric across all settings compared to other methods. For instance, compared to the second-best values, under equispaced 1D subsampling masks of 4x acceleration, DUN-SA improves the PSNR value from SAN’s 46.17db to 47.25db. Meanwhile, under random 1D subsampling masks of 8x acceleration, DUN-SA enhances SAN’s SSIM value from 0.9829 to 0.9846 and reduces the MAE value from 0.0050 to 0.0047.
@@
Fig. 5 illustrates a qualitative comparison of the reconstructed images of different methods and their corresponding error maps on the IXI dataset for 8x acceleration under equispaced 1D subsampling mask. The content displayed is consistent with Fig. 4. Specifically, from the error maps, it’s evident that our method yields the minimal error. Observing the zoomed-in region of interest, in the area indicated by the green arrow, only DUN-SA successfully reconstructs the relevant texture details. In contrast, in the region pointed out by the red arrow, other methods either introduce excessive noise or fail to reconstruct the corresponding texture, but the proposed DUN-SA can reconstruct relatively clear details without being significantly affected by noise. These results suggest that our approach is capable of producing high-quality reconstruction results that present tissues with fewer artifacts and noise compared to other methods.

### v2

@@ 3.3 Network parameters and network training
where $x_{\text{rec},i}$ denotes the $i^{th}$ image reconstructed by the network, and $x_{\text{gt},i}$ signifies the $i^{th}$ ground-truth image.
For $\omega$: we fix $\theta$ and update $\omega$ as follows:
$\omega\leftarrow\omega-\eta_{1}\nabla_{\omega}\mathcal{L}(\omega,\theta^{*})$
where $\eta_{1}$ is the learning rate, and $\nabla_{\omega}\mathcal{L}$ represents the gradient of the loss function with respect to $\omega$.
For $\theta$: While keeping $\omega$ fixed, $\theta$ is updated as follows:
$\theta\leftarrow\theta-\eta_{2}\nabla_{\theta}\mathcal{L}(\omega^{*},\theta)$
where $\eta_{2}$ is the learning rate, and $\nabla_{\omega}\mathcal{L}$ represents the gradient of the loss function with respect to $\theta$.
@@ 4.1 Datasets
We evaluate our method using four datasets, namely the fastMRI dataset, the IXI dataset, an In-house dataset and the BraTS 2018 dataset.
FastMRI Dataset [^1] https://fastMRI.med.nyu.edu/.: To maintain consistency in the experiments, we follow the setups described in [38] and select 340 pairs of T1-weighted and T2-weighted axial brain MRIs. We use 170 volumes (2720 pairs of slices) as the training set, 68 volumes (1088 pairs of slices) as the validation set, and 102 volumes (1632 pairs of slices) as the test set. The in-plane size of all T1-weighted images and T2-weighted images is $320\times 320$, with the resolution of $0.68mm\times 0.68mm$ and the slice spacing of 5mm.
IXI Dataset [^2] http://brain-development.org/ixi-dataset/.: The IXI dataset contains 576 paired multi-modal 3D brain MRIs. We select 570 pairs of PD-weighted and T2-weighted axial brain MRIs for our experiments. Specifically, we use 285 volumes as the training set, 115 volumes as the validation set, and 170 volumes as the test set. The in-plane size of all PD-weighted images and T2-weighted images is $256\times 256$. Consistent with [18], we select the middle 100 slices from each volume for our experiments.
@@
BraTS 2018 Dataset [^3] https://ipp.cbica.upenn.edu/.: The BraTS 2018 dataset includes clinically-acquired, multi-modal MRI scans from patients with glioblastoma, featuring four different imaging modalities: T1, T2, Flair, and T1-weighted contrast-enhanced (T1CE). We selected the high-grade gliomas (HGG) subset, which comprises 210 3D brain MRIs. Specifically, we use 105 volumes as the training set, 42 volumes as the validation set, and 63 volumes as the test set. The in-plane size of all images is $240\times 240$, and we select the middle 20 slices from each volume for our experiments.
@@ 4.2 Compared Methods
Table 1: Quantitative evaluation of DUN-SA vs. other methods on the fastMRI dataset for 4$\times$ and 8$\times$ acceleration under equispaced and random 1D subsampling masks, where T1-weighted images are used as reference modality to assist the reconstruction of T2-weighted images. Best results are emphasized in bold, and the second best are emphasized with an underline.
Methods 4$\times$ Acceleration 8$\times$ Acceleration
@@
Random Zero-filling 27.12$\pm$1.03 0.7373$\pm$0.0282 0.0233$\pm$0.0032 23.88$\pm$1.70 0.5818$\pm$0.0431 0.0342$\pm$0.0067
@@
To evaluate the performance of our model (DUN-SA), we conduct comparisons with state of the art methods, including two single-modal MRI reconstruction methods: E2E-Varnet [28], HQS-Unet [37]; and four multi-modal MRI reconstruction methods: MD-DUN [44], MM-E2E-Varnet(a variant of E2E-VarNet for multi-modal input), SAN [38], and MC-CDic [18]. The details are presented as follows.
E2E-Varnet employs the variational technique to solve the MRI reconstruction problem, unfolding it into a network structure combined with U-Net for end-to-end learning.
4.3 Performance Evaluation Metric
To accurately and comprehensively evaluate the performance of image reconstruction, we select three core metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Mean Absolute Error (MAE). These metrics are employed to quantify the similarity between the reconstructed images and the corresponding ground truth. It is noteworthy that higher values of PSNR and SSIM, along with lower MAE values, indicate superior reconstruction performance of the model.
4.4 Implementation Details
In our experiments, we follow the paradigm set by the fastMRI challenge, where under-sampled MRIs are obtained by masking the corresponding fully-sampled MRIs in k-space using a Cartesian sampling pattern. Specifically, we employ two types of sampling patterns: random and equispaced, to thoroughly assess and compare the performance of DUN-SA with other methods, given their distinct impacts on reconstruction. These are applied at sampling ratios of 25% (4$\times$ acceleration) and 12.5% (8$\times$ acceleration), respectively. Acknowledging that low-frequency signals contain the majority of energy in k-space, we allocate 32% of the sampling to these low frequencies in both patterns. The remaining portion of the sampling is distributed either randomly or in an equispaced manner.
All network models used in these experiments are implemented using the PyTorch framework and are run on a computer equipped with four NVIDIA GeForce GTX 3090 GPUs. During the model training phase, we set the batch size to be $2$, adopt an end-to-end training strategy, and use the SSIM loss function with a coefficient of $1$. We train and test using different slice sizes for each dataset: (320, 320) for the fastMRI and In-house datasets, (256, 256) for the IXI dataset, and (240, 240) for the BraTs 2018 dataset. All parameters are optimized using the Adam optimizer with a learning rate of $1\times 10^{-4}$. To prevent overfitting, we used a validation set to select the optimal parameters during the training process. The final experimental results are based on these optimal parameters, as selected and validated through the validation set, and applied on the test set.
Fig. 4: Visual comparison with representative methods for 4$\times$ acceleration under 1D equispaced subsampling mask on fastMRI dataset. First row: Reconstructed images by different methods; second row: Zoomed-in region of interest; third row: Equispaced mask of 4$\times$ acceleration and error maps of different methods.
@@ 5.1 Results on the fastMRI dataset
Table 2: Quantitative evaluation of DUN-SA vs. other methods on the IXI dataset for 4$\times$ and 8$\times$ acceleration under equispaced and random 1D subsampling masks, where PD-weighted images are used as reference modality to assist the reconstruction of T2-weighted images. Best results are emphasized in bold, and the second best are emphasized with an underline.
Methods 4$\times$ Acceleration 8$\times$ Acceleration
@@
Random Zero-filling 26.32$\pm$2.04 0.6177$\pm$0.0465 0.0296$\pm$0.0068 23.91$\pm$2.01 0.5371$\pm$0.0538 0.0380$\pm$0.0086
@@
Fig. 5: Visual comparison with representative methods for 8$\times$ acceleration under 1D equispaced subsampling mask on the IXI dataset. First row: Reconstructed images by different methods; second row: Zoomed-in region of interest; third row: Equispaced mask of 8$\times$ acceleration and error maps of different methods.
We first run our proposed method on the fastMRI dataset. In this part of the experiments, we use the fully-sampled T1-weighted image as the reference modality to assist the k-space under-sampled data of the T2 modality in reconstruction.
Table 1 displays the quantitative evaluations on the fastMRI dataset for 1D equispaced and random masks under 4× and 8× acceleration. We observe that images reconstructed by multi-modal methods show significant improvements in the PSNR, SSIM, and MAE metrics across all experimental setups compared to those reconstructed by single-modal methods. This enhancement is particularly pronounced for 8$\times$ acceleration. The introduction of spatial alignment in the multi-modal reconstruction further improves the PSNR, SSIM, and MAE values. It is evident that DUN-SA outperforms all others in every metric across all settings. In comparison with the second-best values, for 8$\times$ acceleration under random 1D subsampling masks, the highest differences in PSNR, SSIM, and MAE are 1.38db, 0.0046, and 0.001, respectively. In contrast, for 4$\times$ acceleration under equispaced 1D subsampling masks, the smallest differences are 0.76db, 0.0019, and 0.0002. These results suggest that DUN-SA can effectively utilize the reference modality to aid the MRI reconstruction of the target modality, leading to a noticeable improvement in reconstruction quality.
Table 3: Quantitative evaluation of DUN-SA vs. other methods on the In-house dataset for 4$\times$ and 8$\times$ acceleration under equispaced and random 1D subsampling masks, where T1-weighted images are used as reference modality to assist the reconstruction of T2-weighted images. Best results are emphasized in bold, and the second best are emphasized with an underline.
Methods 4$\times$ Acceleration 8$\times$ Acceleration
@@
Random Zero-filling 26.82$\pm$0.66 0.7075$\pm$0.0144 0.0261$\pm$0.0024 24.27$\pm$0.66 0.5982$\pm$0.0164 0.0359$\pm$0.0033
@@
Fig. 4 illustrates a qualitative comparison of the reconstructed images and their corresponding error maps using various methods on the fastMRI dataset for 4$\times$ acceleration under 1D equispaced mask. The first row displays the reconstruction resluts from different methods, the second row showcases zoomed-in regions of interest, and the third row demonstrates the error maps. The patterns in the error maps represent reconstruction errors; a less intense color on the map (closer to the cool end of the spectrum) signifies superior reconstruction quality. Clearly, the Zero-filling image exhibits aliasing artifacts and lacks anatomical details. Significantly, compared to other methods, our method reconstructs images with minimal visible artifacts and reconstruction discrepancies compared to the ground truth. Specifically, as depicted in the zoomed-in regions of interest shown in the second row, the detail indicated by the green arrow is only recovered by DUN-SA and MC-CDic. Moreover, at the location indicated by the red arrow, MC-CDic reconstruct a texture that originally does not exist, introducing a mistake. These findings verify that our method not only shows remarkable enhancements in all metrics but also reconstructs more textural details.
5.2 Results on the IXI dataset
@@
Fig. 6: Visual comparison with representative methods for 8$\times$ acceleration under 1D equispaced subsampling mask on the In-house dataset. First row: Reconstructed images by different methods; second row: Zoomed-in region of interest; third row: Equispaced mask of 8$\times$ acceleration and error maps of different methods.
Table 4: Quantitative evaluation of DUN-SA vs. other methods on the BraTs 2018 dataset for 4$\times$ and 8$\times$ acceleration under equispaced and random 1D subsampling masks, where the Reference/Target modalities are T2/FLAIR; T1/FLAIR; T1CE/FLAIR; T1/T1CE; T1CE/T2. Best results are emphasized in bold.
Ref/Tgt Methods Equispaced 4$\times$ Equispaced 8$\times$ Random 4$\times$ Random 8$\times$
     T2/FLAIR MC-CDic 44.72 $\pm$ 2.50 37.37 $\pm$ 2.19 42.09 $\pm$ 2.34 37.12 $\pm$ 2.29
0.9906 $\pm$ 0.0034 0.9602 $\pm$ 0.0134 0.9821 $\pm$ 0.0059 0.9567 $\pm$ 0.0148
MM-E2E-Varnet 45.64 $\pm$ 2.54 37.53 $\pm$ 2.25 44.14 $\pm$ 2.45 37.84 $\pm$ 2.33
0.9920 $\pm$ 0.0031 0.9603 $\pm$ 0.0128 0.9886 $\pm$ 0.0042 0.9613 $\pm$ 0.0131
SAN 46.03 $\pm$ 2.63 38.41 $\pm$ 2.35 44.60 $\pm$ 2.48 37.88 $\pm$ 2.31
0.9926 $\pm$ 0.0029 0.9645 $\pm$ 0.0125 0.9895 $\pm$ 0.0037 0.9613 $\pm$ 0.0131
DUN-SA 48.14 $\pm$ 2.83 39.43 $\pm$ 2.40 46.93 $\pm$ 2.66 39.12 $\pm$ 2.42
0.9950 $\pm$ 0.0023 0.9694 $\pm$ 0.0111 0.9932 $\pm$ 0.0027 0.9682 $\pm$ 0.0114
     T1/FLAIR MC-CDic 44.83 $\pm$ 2.43 37.40 $\pm$ 2.03 41.86 $\pm$ 2.14 37.10 $\pm$ 2.06
0.9907 $\pm$ 0.0032 0.9598 $\pm$ 0.0112 0.9815 $\pm$ 0.0059 0.9567 $\pm$ 0.0120
MM-E2E-Varnet 46.03 $\pm$ 2.53 38.05 $\pm$ 2.02 43.82 $\pm$ 2.19 37.66 $\pm$ 2.11
0.9927 $\pm$ 0.0028 0.9627 $\pm$ 0.0126 0.9878 $\pm$ 0.0078 0.9603 $\pm$ 0.0110
SAN 46.03 $\pm$ 2.63 38.09 $\pm$ 2.03 44.60 $\pm$ 2.21 37.67 $\pm$ 2.10
0.9926 $\pm$ 0.0029 0.9627 $\pm$ 0.0134 0.9895 $\pm$ 0.0081 0.9603 $\pm$ 0.0109
DUN-SA 48.34 $\pm$ 2.81 39.03 $\pm$ 2.19 46.51 $\pm$ 2.57 38.81 $\pm$ 2.24
0.9952 $\pm$ 0.0021 0.9677 $\pm$ 0.0097 0.9927 $\pm$ 0.0027 0.9666 $\pm$ 0.0102
     T1CE/FLAIR MC-CDic 43.17 $\pm$ 2.24 37.31 $\pm$ 2.28 40.96 $\pm$ 2.26 36.94 $\pm$ 2.26
0.9877 $\pm$ 0.0038 0.9582 $\pm$ 0.0129 0.9789 $\pm$ 0.0136 0.9552 $\pm$ 0.0136
MM-E2E-Varnet 46.22 $\pm$ 2.62 37.87 $\pm$ 2.15 44.16 $\pm$ 2.44 37.53 $\pm$ 2.15
0.9930 $\pm$ 0.0028 0.9611 $\pm$ 0.0149 0.9887 $\pm$ 0.0037 0.9587 $\pm$ 0.0120
SAN 46.37 $\pm$ 2.63 38.02 $\pm$ 2.27 44.04 $\pm$ 2.42 37.60 $\pm$ 2.25
0.9932 $\pm$ 0.0028 0.9618 $\pm$ 0.0114 0.9882 $\pm$ 0.0040 0.9593 $\pm$ 0.0119
DUN-SA 48.19 $\pm$ 2.84 38.98 $\pm$ 2.32 46.50 $\pm$ 2.65 38.70 $\pm$ 2.38
0.9951 $\pm$ 0.0022 0.9670 $\pm$ 0.0104 0.9926 $\pm$ 0.0028 0.9655 $\pm$ 0.0110
     T1/T1CE MC-CDic 48.60 $\pm$ 2.52 41.71 $\pm$ 2.39 46.10 $\pm$ 2.40 41.51 $\pm$ 2.35
0.9935 $\pm$ 0.0029 0.9764 $\pm$ 0.0109 0.9894 $\pm$ 0.0047 0.9760 $\pm$ 0.0110
MM-E2E-Varnet 50.18 $\pm$ 1.98 41.51 $\pm$ 2.27 48.47 $\pm$ 2.32 42.05 $\pm$ 2.44
0.9959 $\pm$ 0.0023 0.9769 $\pm$ 0.0104 0.9933 $\pm$ 0.0033 0.9778 $\pm$ 0.0103
SAN 50.22 $\pm$ 2.66 42.24 $\pm$ 2.41 48.70 $\pm$ 2.62 42.13 $\pm$ 2.44
0.9960 $\pm$ 0.0022 0.9788 $\pm$ 0.0101 0.9936 $\pm$ 0.0043 0.9782 $\pm$ 0.0102
DUN-SA 52.25 $\pm$ 3.05 43.47 $\pm$ 2.59 51.35 $\pm$ 2.97 43.22 $\pm$ 2.60
0.9971 $\pm$ 0.0018 0.9820 $\pm$ 0.0093 0.9960 $\pm$ 0.0023 0.9809 $\pm$ 0.0095
     T1CE/T2 MC-CDic 44.19 $\pm$ 1.77 37.29 $\pm$ 2.17 41.36 $\pm$ 1.69 36.17 $\pm$ 1.83
0.9930 $\pm$ 0.0023 0.9711 $\pm$ 0.0116 0.9856 $\pm$ 0.0044 0.9651 $\pm$ 0.0115
MM-E2E-Varnet 46.61 $\pm$ 1.99 37.43 $\pm$ 1.84 44.54 $\pm$ 1.84 37.32 $\pm$ 1.92
0.9958 $\pm$ 0.0016 0.9713 $\pm$ 0.0130 0.9921 $\pm$ 0.0026 0.9709 $\pm$ 0.0104
SAN 46.73 $\pm$ 2.00 37.73 $\pm$ 1.86 44.78 $\pm$ 1.88 37.38 $\pm$ 1.86
0.9959 $\pm$ 0.0015 0.9726 $\pm$ 0.0101 0.9924 $\pm$ 0.0026 0.9711 $\pm$ 0.0104
DUN-SA 49.22 $\pm$ 2.22 39.29 $\pm$ 2.40 48.07 $\pm$ 2.37 39.03 $\pm$ 1.98
0.9974 $\pm$ 0.0012 0.9782 $\pm$ 0.0073 0.9958 $\pm$ 0.0017 0.9774 $\pm$ 0.0092
Fig. 7: Visual comparison with representative methods for 8$\times$ acceleration under 1D random subsampling mask on the BraTs 2018 dataset. First row: Reconstructed images by different methods; second row: Zoomed-in region of interest; third row: Random mask of 8$\times$ acceleration and error maps of different methods.
@@
We validate the effectiveness of our method on a larger dataset, IXI. In this part of the experiments, the fully sampled PD-weighted image is used as reference modality image to assist the reconstruction of the paired under-sampled T2-weighted image.
@@
Table 2 presents the quantitative evaluations on the IXI dataset for 1D equispaced and random masks under 4× and 8× acceleration. From the metrics, it can be observed that the reference modality can significantly enhance the reconstruction performance, and the spatial alignment further improves all metrics, which is consistent with the experimental results on the fastMRI dataset. Our method once again brings noticeable improvements for every metric across all settings compared to other methods. For instance, compared to the second-best values, under equispaced 1D subsampling masks of 4$\times$ acceleration, DUN-SA improves the PSNR value from SAN’s 46.17db to 47.25db. Meanwhile, under random 1D subsampling masks of 8$\times$ acceleration, DUN-SA enhances SAN’s SSIM value from 0.9829 to 0.9846 and reduces the MAE value from 0.0050 to 0.0047.
@@
Fig. 5 illustrates a qualitative comparison of the reconstructed images of different methods and their corresponding error maps on the IXI dataset for 8$\times$ acceleration under equispaced 1D subsampling mask. The content displayed is consistent with Fig. 4. Specifically, from the error maps, it’s evident that our method yields the minimal error. Observing the zoomed-in region of interest, in the area indicated by the green arrow, only DUN-SA successfully reconstructs the relevant texture details. In contrast, in the region pointed out by the red arrow, other methods either introduce excessive noise or fail to reconstruct the corresponding texture, but the proposed DUN-SA can reconstruct relatively clear details without being significantly affected by noise. These results suggest that our approach is capable of producing high-quality reconstruction results that present tissues with fewer artifacts and noise compared to other methods.